OpenStudy (happpily):

Use the quadratic formula to solve the equation 4x^2 - 6x 10 = 0

OpenStudy (asett):

Is there something missing between the 6x and 10?

OpenStudy (happpily):

yeah im so sorry its \[4^{2}- 6x - 10 = 0\]

OpenStudy (asett):

np. Do you see anything we can take out?

OpenStudy (asett):

We can take out a 2: from 4x^2, 6x, and 10, leaving us with 2x^2, 3x, and 5.

OpenStudy (happpily):

oh ok got it.

OpenStudy (asett):

Yay! So now, we are left with\[2(2x ^{2}-3x-5)\]

OpenStudy (asett):

The cool thing about this is, do you notice that the difference between the coefficient of 2x^2 and 5 is -3?

OpenStudy (happpily):

oh true

OpenStudy (asett):

So, my math teacher told me this (it is so cool!). When we split it into two groups, the 2x and the -5 have to be together! Based on this, can you come up with the answer?

OpenStudy (happpily):

i got x = 5/2 x= 1 .... ?

OpenStudy (anonymous):

the formula given is wrong here lol

OpenStudy (asett):

Are you sure that it is +1?

OpenStudy (asett):

@ChiefArnav ... techinical issues (like my spelling)

OpenStudy (happpily):

oh no its negative!!

OpenStudy (happpily):

sorry about that!

OpenStudy (asett):

Correct! Great job. Movin' right along...

OpenStudy (asett):

No problem, I have a hard time keeping my positives/negatives straight too. : )

OpenStudy (happpily):

thank you for helping me figure it out once again! :D

OpenStudy (asett):

Your welcome!